OpenStudy (anonymous):

how can properties of logarithms be applied to simplify or solve equations??????

OpenStudy (anonymous):

take an exponential equation. i.e x equals something to the power of x. We can use a logarithm to solve this. i.e. to find that x

OpenStudy (anonymous):

They can be used to combine multiple log expressions into one. For example \[\log (x+3)-\log(x+5)=10\] \[\log \frac{ x+3 }{ x-5 }=10\] Then you can convert that to an exponential equation and solve fairly easily

OpenStudy (anonymous):

Thank you so much! How do I convert it to exponential??

OpenStudy (anonymous):

You need to use @beezle1197 to get beezle1197